I need help on solving a proportion problem. Could you help me by showing me step by step. 3r-2 all over r = 7 over 8
Answer by Theo(13342)

you have:
(3r-2) / r = 7/8
multiply both sides of this equation by r to get:
(3r-2) = r * (7/8)
multiply both sides of this equation by 8 to get:
8 * (3r - 2) = r * 7
simplify this equation to get:
24r - 16 = 7r
add 16 to both sides of this equation and subtract 7r from both sides of this equation to get:
24r - 7r = 16
simplify to get:
17r = 16
divide both sides of this equation by 17 to get:
r = 16/17
substitute 16/17 for r in your original equation to see if the equation holds true.

your original equation is:
(3r-2) / r = 7/8
substitute for r to get:
(3 * (16/17) - 2) / (16/17) = 7/8
multiply both sides of this equation by (16/17) to get:
(3 * (16/17) - 2) = (7/8) * (16/17)
simplify this equation to get:
(48/17 - 2) = (112/136)
simplify further to get:
(48/17 - 34/17) = (112/136)
simplify further to get:
(14/17) = (112/136)
112/136 is equivalent to 14/17 so you get:
(14/17) = (14/17) which is true, confirming the value of (16/17) for r is good.